Find A so that the point (1,1) is on the graph of f(x)=axsquared+4
Answer by Alan3354(69443)

Find A so that the point (1,1) is on the graph of f(x)=axsquared+4
------
Assuming A = a:
f(x) = ax^2 + 4
f(1) = a + 4 = 1
a = -3